OpenStudy (anonymous):

Tickets for the regional volleyball tournament went on sale at Lazy Creek High School on Monday. General admission tickets cost $6.00 each and reserved seating tickets cost $12.00 each. If 400 tickets were sold on Monday and the total ticket sales was $3150.00, how many of each type of ticket was sold? Question 6 options: a) general admission tickets=275; reserved seating tickets=125 b) general admission tickets=125; reserved seating tickets=275 c) general admission tickets=225; reserved seating tickets=175 d) none of these

OpenStudy (anonymous):

so can u figure out the two equations you need?

OpenStudy (anonymous):

what types of tickets are there?

OpenStudy (anonymous):

General admission and reserved

OpenStudy (anonymous):

ok lets let General admission be G and reserved be R. it says the total number of tickets sold was 400 can you model that in one equation?

OpenStudy (anonymous):

g+r = 400???

OpenStudy (anonymous):

good. now if general admission costs 6 bucks and reserved cost 12 bucks, can you model the formula to find out how much you make from selling g and d amounts of tickets?

OpenStudy (anonymous):

6g + 12r = 3150.00????

OpenStudy (anonymous):

see its not that hard :)

OpenStudy (anonymous):

now solve the system together. i recommend substitution method since you have variables without numeric coefficients (numbers in front of them)

OpenStudy (anonymous):

can you solve G+R=400 for G?

OpenStudy (anonymous):

g = -r + 400

OpenStudy (anonymous):

ok now plug that G into 6G+12R=3150

OpenStudy (anonymous):

6(-r + 400) + 12r = 31500

OpenStudy (anonymous):

solve for R

OpenStudy (anonymous):

notice in your problem the answers all have different Rs (hint hint)

OpenStudy (anonymous):

-6r + 2400 + 12r = 31500 6r + 2400 = 31500 6r = 29100 r = 4850 Uh. What did I do wrong???

OpenStudy (anonymous):

its 3150 not 31500

OpenStudy (anonymous):

....oh.

OpenStudy (anonymous):

oops i should of caught that above my bad

OpenStudy (anonymous):

r = 125

OpenStudy (anonymous):

so what's the answer above?

OpenStudy (anonymous):

wait. r = 750 doesn't make sense either....

OpenStudy (anonymous):

no R = 125

OpenStudy (anonymous):

G = 400-R what's G

OpenStudy (anonymous):

I forgot to write down the divided part XDDD g = 275

OpenStudy (anonymous):

ok which answer above has G=275 and R= 125?

OpenStudy (anonymous):

a!!!

OpenStudy (anonymous):

:)

OpenStudy (texaschic101):

good explaining druminjosh :)

OpenStudy (anonymous):

thanks @texaschic101

OpenStudy (anonymous):

its because i'm from Texas too :)

OpenStudy (anonymous):

XD